OpenStudy (anonymous):

solve for variable: x^4-10x^2+9=0

OpenStudy (anonymous):

Simplifying x4 + -10x2 + 9 = 0 Reorder the terms: 9 + -10x2 + x4 = 0 Solving 9 + -10x2 + x4 = 0 Solving for variable 'x'. Factor a trinomial. (1 + -1x2)(9 + -1x2) = 0 Factor a difference between two squares. ((1 + x)(1 + -1x))(9 + -1x2) = 0 Factor a difference between two squares. ((3 + x)(3 + -1x))(1 + x)(1 + -1x) = 0 Subproblem 1Set the factor '(1 + x)' equal to zero and attempt to solve: Simplifying 1 + x = 0 Solving 1 + x = 0 Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + x = 0 + -1 Combine like terms: 1 + -1 = 0 0 + x = 0 + -1 x = 0 + -1 Combine like terms: 0 + -1 = -1 x = -1 Simplifying x = -1 Subproblem 2Set the factor '(1 + -1x)' equal to zero and attempt to solve: Simplifying 1 + -1x = 0 Solving 1 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-1' to each side of the equation. 1 + -1 + -1x = 0 + -1 Combine like terms: 1 + -1 = 0 0 + -1x = 0 + -1 -1x = 0 + -1 Combine like terms: 0 + -1 = -1 -1x = -1 Divide each side by '-1'. x = 1 Simplifying x = 1 Subproblem 3Set the factor '(3 + x)' equal to zero and attempt to solve: Simplifying 3 + x = 0 Solving 3 + x = 0 Move all terms containing x to the left, all other terms to the right. Add '-3' to each side of the equation. 3 + -3 + x = 0 + -3 Combine like terms: 3 + -3 = 0 0 + x = 0 + -3 x = 0 + -3 Combine like terms: 0 + -3 = -3 x = -3 Simplifying x = -3 Subproblem 4Set the factor '(3 + -1x)' equal to zero and attempt to solve Simplifying 3 + -1x = 0 Solvinh 3 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-3' to each side of the equation. 3 + -3 + -1x = 0 + -3 Combine like terms: 3 + -3 = 0 0 + -1x = 0 + -3 -1x = 0 + -3 Combine like terms: 0 + -3 = -3 -1x = -3 Divide each side by '-1'. x = 3 Simplifying x = 3Solutionx = {-1, 1, -3, 3}

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

yess ! :D THANK YOU!

OpenStudy (anonymous):

yo welcome